OpenStudy (diyadiya):

Discuss the continuity of the cosine, cosecant, secant and cotangent functions,

OpenStudy (anonymous):

cosine is always continous

OpenStudy (mr.math):

Yep!

OpenStudy (anonymous):

cosecant, secant and cotangent are discontinous at a point where their inverse are 0

OpenStudy (mr.math):

\(\csc(x)=\frac{1}{\sin(x)}\), discontinuous at zeros of \(\sin(x)\).

OpenStudy (mr.math):

Same thing with secant, discontinuous at zeros of \(\cos(x)\).

OpenStudy (diyadiya):

how should i write it ?

OpenStudy (mr.math):

I will help you with co-secant: Let \(f(x)=\csc(x)=\frac{1}{\sin(x)}\). The zeros of \(\sin(x)\) are \(\frac{\pi}{2}+2n\pi\), where n is any integer. Thus f(x) is discontinuous at \(\frac{\pi}{2}+2n\pi\), \(n\in \Integer\).

OpenStudy (anonymous):

Cosine is continuous at all points. Secant is discontinuous at the points where the value of the variant is odd multiple of pi/2 Cosecant is discontinuous at the points where the variant is a multiple of pi Cotangent is discontinuous at the points where the variant is a multiple of pi.

OpenStudy (mr.math):

Crap! Zeros of \(\sin(x)\) are multiples of \(2\pi\).

OpenStudy (mr.math):

So cosecant is discontinuous at \(x=2n\pi\), where n is any integer.

OpenStudy (diyadiya):

Yeah ! :)

OpenStudy (mr.math):

:D

OpenStudy (mr.math):

What I did above is the zeros of cos(x), which states the discontinuity of \(\sec(x)\).

OpenStudy (zarkon):

isn't it just the multiples of \(\pi\)

OpenStudy (zarkon):

not \(2\pi\)

OpenStudy (mr.math):

Lol, what's wrong with me?!

OpenStudy (diyadiya):

is it?

OpenStudy (mr.math):

Yeah, as Zarkon said zeros of \(\sin(x)\) are \(n\pi\), and zeros of \(\cos(x)\) are \(\frac{\pi}{2}+n\pi\).

OpenStudy (mr.math):

Sorry for the confusion!

OpenStudy (mr.math):

And thanks for Zarkon!

OpenStudy (diyadiya):

And Tan ?

OpenStudy (mr.math):

What do you think?

OpenStudy (diyadiya):

hm lol idk :D

OpenStudy (mr.math):

Use the same concept (i.e find zeros of the denominator).

OpenStudy (mr.math):

\(\tan(x)=\frac{\sin(x)}{\cos(x)}\), so?

OpenStudy (mr.math):

It would have the same discontinuity points as \(\sec(x)\).

OpenStudy (mr.math):

While \(\cot(x)\) will have the same discontinuity points as?

OpenStudy (diyadiya):

So cos x?

OpenStudy (diyadiya):

i mean discontinouos at zeroes of cos x?

OpenStudy (mr.math):

Yeah, tan(x) is discontinouos at zeros of \(\cos(x)\), true!

OpenStudy (mr.math):

Try to reread the last few statements I made.

OpenStudy (diyadiya):

Sure :) ThankYOu Libniz ,Mr.Math,Atchyut & Zarkon :)

OpenStudy (mr.math):

Good luck Diya! :D

OpenStudy (diyadiya):

Thanks :)